OpenStudy (anonymous):

Q: Given that n varies directly with m with the constant k = 5/6 if n=9 find m to the nearest tenth A. m=7.5 B. m=10.8 C. m=-7.5 D. m=6.4

OpenStudy (abmon98):

n=km 9=5/6m m=6*9/5